OpenStudy (anonymous):

Complete the square to solve the equation. x^2-10x-2=17

OpenStudy (anonymous):

start with \[x^2-10x=19\] half of 10 is 5, 5 squared is 25 so add 25 to get \[(x-5)^2=19+25\] \[(x-5)^2=44\] \[x-5=\pm\sqrt{44}\] \[x=5\pm\sqrt{44}\] \[x=5\pm2\sqrt{11}\]

OpenStudy (anonymous):

so im confused what is the answer

OpenStudy (anonymous):

which step was not clear?

OpenStudy (anonymous):

heromiles go away you're annoying and your stupid little comments don't help.

OpenStudy (anonymous):

What did I do?

OpenStudy (anonymous):

I'm appalled :/

OpenStudy (anonymous):

we can work a different one out step by step if you like

OpenStudy (anonymous):

say \[x^2+6x-2=0\]

OpenStudy (anonymous):

It seems like you always have a rude comment to say so if you're not going to help then flutter off

OpenStudy (anonymous):

I didn't make a rude comment. I was simply letting you know that satellite did indeed post the answer. Where was I rude?????? :/

OpenStudy (anonymous):

& I get it satellite I just overlooked it thank you

OpenStudy (anonymous):

I don't recall making a "rude" comment as you say....:/

OpenStudy (anonymous):

I was trying to help

OpenStudy (anonymous):

well it seemed like you were being rude im not really about to discuss this with you. your way of "trying to help" isn't the helpful its pretty mean the way i see it.

OpenStudy (anonymous):

Hmmm, you asked "What's the answer?". I simply replied, "Satellite posted the answer already". I wasn't trying to be mean at all....I promise....

OpenStudy (anonymous):

I also helped you with a couple questions before...it's just that satellite posted the answers before I did. I would have explained to you exactly how to solve them.

OpenStudy (anonymous):

you said "hmm seems like he posted it already?" which seemed like it was sacastic and you were trying to undermine my intelligence and i don't have time for that.

OpenStudy (anonymous):

I'm sorry if you think I was being rude...I don't have time for that either. I don't like when people are rude to me. I would never.....

OpenStudy (anonymous):

I deleted the rude comment :/

OpenStudy (anonymous):

It's fine. All is forgiven.